This needs to be solved by the substitution method.
4x + 4y = 0
-8x + y = 72

Answer by mananth(16946)

4x+4y=0 .............1
y=72+8x ---------2
substitute y in (1)
4x+4(72+8x)=0
4x+288+32x=0
4x+32x=0-288
36x=-288
/36
x=-8
Plug the value of x in (1)
4x+4y=0
4*-8+4y=0
-32+4y=0
4y=32
/4
y=8